# An Investigation into the Distribution of Ratios of Particle Solver-based Likelihoods

### v1

@@ 1 Introduction
$\partial_{t}\rho(x,t)=D\partial_{xx}\rho(x,t),$ (1)
$\rho(0,t)=\rho(L,t),\qquad\partial_{x}\rho(0,t)=\partial_{x}\rho(L,t),\qquad\rho(x,0)=\rho_{0}(x)\geq 0.$ (2)
Here, $t\in\mathbb{R}_{\geq 0}$ is the time variable and $\rho_{0}(x)$ is a given initial condition. We approximate the solution to this equation using a Monte Carlo simulation. Such simulations discretize the equation at hand using an ensemble of particles. The particle dynamics are chosen such that the ensemble statistics converge to a (potentially biased) approximation of the true solution of (1)–(2) as the particle ensemble size $P\to\infty$. Although Monte Carlo simulation can not compete with deterministic approaches, such as finite differences [12] or spectral methods [11] for the problem at hand, we are motivated by their indispensable character in various high-impact domains, including nuclear fusion research [6, 20], financial mathematics [7] and computer graphics [19], considering the diffusion equation as a straightforward, low-cost example.
@@
where, $G[\cdot]:\mathbb{R}\mapsto\mathbb{R}^{N}$ is the solution map of (1)–(2) for a given diffusion coefficient $D\in\mathbb{R}$ on a discrete spatial grid. The observation error $\eta\in\mathbb{R}^{N}$ is an unknown multivariate random quantity, assumed normally distributed with a strictly positive definite covariance matrix $\Sigma_{\eta}=\Sigma_{\eta}^{\mathsf{T}}\!\in\mathbb{R}_{\succ 0}^{N\times N}$. Although $\eta$ is modeled as a random variable, it takes a fixed, but unknown, value for a given observation $\rho$.
@@
resulting from the use of Monte Carlo simulation. Here, $\delta\in\mathbb{R}^{N}$ follows a normal distribution with mean $\mu:\mathbb{R}\mapsto\mathbb{R}^{N}$ and covariance matrix $\Sigma=\Sigma^{\mathsf{T}}\!:\mathbb{R}\mapsto\mathbb{R}^{N\times N}_{\succ 0}$, parameterized by the parameter $D$. We motivate the assumed distribution in Section 2. As $\hat{G}$$\cdot$] is not easily invertible, solving the inverse problem (3) requires multiple evaluations of (1)–(2[23]. This fact, combined with the high computational cost of Monte Carlo simulation means a cost-accuracy trade-off must be made in the value of $P$. This trade-off is further exacerbated when multiple evaluations of the model are needed, such as when solving problems of the form (3). One must then take into account both the stochastic nature of the approximation error, as well as the generally high simulation cost incurred.
Due to this computational cost, most efforts to solve inverse problems using Monte Carlo simulation have relied on neglecting the observation error $\eta$, to solve the inverse problem as an optimization problem, minimizing the residual G[D]yG[D]-y under a suitable norm. The main challenge in the setting given by (4), where the forward model contains a solver with stochastic errors, is computing gradients. Various approaches have been applied to tackle this challenge, notably adjoint methods, see e.g. [4, 15, 13, 2, 3], and algorithmic differentiation, see e.g. [10, 25, 26]. However, in this work, we consider a Bayesian approach, where we approximate a posterior distribution of $D$ conditioned on the observation $\rho$. The Bayesian posterior is viewed as the solution to the inverse problem as it encodes all of the available information about the unknown $D$ given the data $\rho$.
@@
[…]
In our setting, the posterior given by (5) will depend nonlinearly on the random error $\delta$, due to using the approximate solution map (4) in the log-likelihood. Given a deterministic discretization error $\delta$, one can apply standard textbook results to bound the resulting error on the posterior (5) in, e.g., Hellinger distance [22, Ch. 1]. Similar results on posterior convergence were proven in [14] for random errors $\delta$, under suitable conditions, for the limit $\delta\to 0$. However, we consider the case of non-negligible stochastic errors. To the best of our knowledge, no theoretical results on well-posedness exist in this setting. We focus on algorithmic analysis in this work, rather than the posterior itself.
@@
In the considered setting with stochastic errors (4), we note that the use of MCMC for sampling the posterior (5) results in a nested sampling problem as each MCMC sample requires a full Monte Carlo simulation. Our interest lies in understanding the effect of non-negligible noise in the forward map. For inherently random but unbiased likelihoods, one can rely on pseudo-marginal MCMC methods [1]. Such likelihoods are often the result of a product or sum of individual contributions from different data points [16] or simulation outputs [24]. Pseudo-marginal methods consider the combined probability space of the parameter to be inferred and the stochastic dimension of the forward map. The aim is to approximate the marginal distribution of the inferred parameter $D$, i.e., integrating out the forward map’s stochasticity, using samples from this combined probability space. Although these methods are well studied, the assumption of unbiasedness means that they cannot be directly applied in our setting.
The goal of this work is to consider the non-asymptotic sampling regime in $P$, i.e., simulations with non-negligible stochastic error. We study how the Monte Carlo simulation error affects the ratio of computed approximate likelihood values, as a first step towards understanding the acceptance-rejection behavior of the Metropolis-Hastings algorithm in this setting. Notably, we observe that this ratio becomes a random variable, whose distribution we study in terms of its moments. This line of work complements that presented in [17], applying MCMC to Monte Carlo simulations in photoacoustic imaging.
@@ 2 Monte Carlo simulation of the diffusion equation
In this Section, we derive a Monte Carlo discretization of the model (1)–(2) and comment on the structure of the resulting error. This derivation is purposefully brief. For a more detailed mathematical derivation of Monte Carlo simulation for parabolic PDEs, we refer to textbooks, such as [8]. We assume, for simplicity, that $\rho(x,t)\geq 0$. As such, we can interpret $\rho(x,t)$ as a probability density over $x$ at any given time $t$. We then make use of the fact that (1) is the Fokker-Planck equation to the stochastic differential equation (SDE)
@@
at time $t=k\Delta t$. Using an Euler-Maruyama discretization, we propagate each particle $p$ in (7) as a discretized realization of (6), i.e., for $p=1,\dots,P$
@@
and $X_{p,0}\overset{\text{i.i.d.}}{\sim}\rho_{0}$. Simulating such an ensemble, produces an ensemble of particles, where the values $X_{p,k}$ are distributed according to $\rho(x,t)$ for any $t=k\Delta t$.
@@
We define the error $\delta_{k}\in\mathbb{R}^{N}$ of the Monte Carlo simulation, at time $t=k\Delta t$, using the elementwise notation
$\delta_{n,k}=\rho(x_{n},t)-\hat{\rho}_{n,k},\qquad n=0,\dots,N-1,\qquad k\geq 0.$ (10)
@@
$\mathbb{E}\left[\hat{\rho}_{n,k}\right]=\frac{1}{\Delta x}\int_{n\Delta x}^{(n+1)\Delta x}\rho(x,t)\text{d}x=\frac{F((n+1)\Delta x,t)-F(n\Delta x,t)}{\Delta x},$ (12)
@@
$\mathbb{E}\left[\hat{\rho}_{n,k}\right]=\rho(x,t)+\left(\left(n+\frac{1}{2}\right)\Delta x-x\right)\frac{\partial\rho}{\partial x}(x,t)+\mathcal{O}(\Delta x^{2}).$ (13)
@@
$\left(n+\frac{1}{2}\right)\Delta x-x\leq\frac{\Delta x}{2}.$ (14)
@@
The finite ensemble size $P$ induces a sampling error that scales with the number of particles in a given cell. As each particle in the ensemble (7) is fully independent and produces a single contribution to the binning estimator (9), we observe through the law of large numbers that the variance of the individual grid cells scales with $\mathcal{O}\left(\frac{1}{P\Delta x}\right)$ as $P\Delta x\to\infty$. Moreover, the distribution of $\delta_{k}$ will converge to a multivariate Gaussian in this limit, with mean and variance corresponding to the bias and variance in the error decomposition (11). Hence, motivating the assumed distribution in (4). Despite the trajectory independence, we note that the grid cells themselves are correlated due to the fact that each trajectory can only score in one cell at a given time step.
@@ 3 Analysis of approximate likelihood ratios
is a random variable, who’s $p$-th moment exists iff $p\Sigma_{\eta}\succ\Sigma(D_{2})$, with $\succ$ denoting the Löwner order.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 Numerical experiments
We now perform a numerical experiment to confirm the observations in Theorem 3.1. To this end, we use the model problem (1)–(2) and consider transitions between fixed values for $D_{1}$ and $D_{2}$. We produce synthetic observations, using a finite difference discretization whose solution is perturbed with noise distributed according to (3). Throughout this section we set $N=100$, $L=10$, $t=10$ and $D^{\ast}=0.1$. The reference solution to produce the observation is computed with $\Delta t=0.1$, while the Monte Carlo simulations use $\Delta t=t$ to cut on computational cost, thanks to the unbiasedness of (8). For $D_{1}$ and $D_{2}$, we consider differing combinations of these parameters from the set $\{0.08,0.1\}$, i.e., the true value and a similar, but less probable, value. The code used to generate the results in this section can be found at github.com/UQatKIT/FrontUQ-2024-proceedings-likelihood-ratios.
@@
Figure 1: Computing the expectation of the approximate likelihoods and the ratio of these expectations, given a discrete solution of the problem (1)–(2) perturbed by synthetic observation noise, for $D^{\ast}=D_{1}=0.1$ and $D_{2}=0.08$.
@@
Figure 2: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (1)–(2) perturbed by synthetic observation noise for $D^{\ast}=D_{1}=D_{2}=0.1$.
@@
Figure 3: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (1)–(2) perturbed by synthetic observation noise for $D^{\ast}=D_{2}=0.1$ and $D_{1}=0.08$.
@@
Figure 4: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (1)–(2) perturbed by synthetic observation noise for $D^{\ast}=D_{1}=0.1$ and $D_{2}=0.08$.
Next, in Figure 2, we consider a transition from $D_{2}=0.1$ to $D_{1}=0.1$. We plot the expectation of the approximate likelihood ratio in Figure 2(a). Here we observe a significant blowup in the expectation at approximately $\sigma_{\eta}=\sigma_{\delta}$, as predicted by Theorem 3.1. For sufficiently small values of $\sigma_{\delta}$, we observe that the expectation of the likelihood ratio converges to the true value of 1. To better understand the divergence of the approximate likelihood ratio, we plot the expectation of the minimum of the ratio and 1 in Figure 2(b), which in the case of $D_{1}=D_{2}$ directly corresponds to the acceptance probability in the Metropolis-Hastings algorithm, see e.g. [21, Ch. 7]. Here, it becomes clear that when the approximate likelihood ratio diverges, the acceptance probability converges to the value 0.5. This behavior gives insight into the precise mechanism behind the divergence. Namely, in this regime, the stochastic error in the approximate likelihood evaluations is sufficiently large that the ratio consists of a numerator and denominator with wildly differing orders of magnitude. Hence, the random variable converges to a distribution with two states, one of which diverges to infinity, the other converges to zero. Each state has probability 0.5.

### v2

@@ 1 Introduction
$\ignorespaces\ignorespaces\partial_{t}\rho(x,t)=D\partial_{xx}\rho(x,t),$
$\rho(0,t)=\rho(L,t),\qquad\partial_{x}\rho(0,t)=\partial_{x}\rho(L,t),\qquad\rho(x,0)=\rho_{0}(x)\geq 0.\ignorespaces\ignorespaces$
Here, $t\in\mathbb{R}_{\geq 0}$ is the time variable and $\rho_{0}(x)$ is a given initial condition. We approximate the solution to this equation using a Monte Carlo simulation. Such simulations discretize the equation at hand using an ensemble of particles. The particle dynamics are chosen such that the ensemble statistics converge to a (potentially biased) approximation of the true solution of (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) as the particle ensemble size $P\to\infty$. Although Monte Carlo simulation can not compete with deterministic approaches, such as finite differences [15] or spectral methods [14] for the problem at hand, we are motivated by their indispensable character in various high-impact domains, including nuclear fusion research [9, 24], financial mathematics [10] and computer graphics [22], considering the diffusion equation as a straightforward, low-cost example.
@@
where, $G[\cdot]:\mathbb{R}\mapsto\mathbb{R}^{N}$ is the solution map of (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) for a given diffusion coefficient $D\in\mathbb{R}$ on a discrete spatial grid. The observation error $\eta\in\mathbb{R}^{N}$ is an unknown multivariate random quantity, assumed normally distributed with a strictly positive definite covariance matrix $\Sigma_{\eta}=\Sigma_{\eta}^{\mathsf{T}}\!\in\mathbb{R}_{\succ 0}^{N\times N}$. Although $\eta$ is modeled as a random variable, it takes a fixed, but unknown, value for a given observation $\rho$.
@@
resulting from the use of Monte Carlo simulation. Here, $\delta\in\mathbb{R}^{N}$ follows a normal distribution with mean $\mu:\mathbb{R}\mapsto\mathbb{R}^{N}$ and covariance matrix $\Sigma=\Sigma^{\mathsf{T}}\!:\mathbb{R}\mapsto\mathbb{R}^{N\times N}_{\succ 0}$, parameterized by the parameter $D$. In other words, $\delta$ is modeled as a white-noise Gaussian random process [23] with respect to $D$. We motivate the assumed distribution in Section 2.
As $\hat{G}$$\cdot$] is not easily invertible, solving the inverse problem (2) requires multiple evaluations of (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary[27]. This fact, combined with the high computational cost of Monte Carlo simulation means a cost-accuracy trade-off must be made in the value of $P$. This trade-off is further exacerbated when multiple evaluations of the model are needed, such as when solving problems of the form (2). One must then take into account both the stochastic nature of the approximation error, as well as the generally high simulation cost incurred.
Due to this computational cost, most efforts to solve inverse problems using Monte Carlo simulation have relied on neglecting the observation error $\eta$, to solve the inverse problem as an optimization problem, minimizing the residual $G[D]-\rho$ under a suitable norm. The main challenge in the setting given by (3), where the forward model contains a solver with stochastic errors, is computing gradients. Various approaches have been applied to tackle this challenge, notably adjoint methods, see e.g. [6, 18, 16, 3, 4], and algorithmic differentiation, see e.g. [13, 29, 30]. However, in this work, we consider a Bayesian approach, where we approximate a posterior distribution of $D$ conditioned on the observation $\rho$. The Bayesian posterior is viewed as the solution to the inverse problem as it encodes all of the available information about the unknown $D$ given the data $\rho$.
@@
where $\pi_{0}(\cdot)$ is the PDF of an assumed prior distribution on $D$, $\pi_{\text{l}}(\rho|D)$ is the likelihood of $\rho$ given $D$, and $\pi_{\text{obs}}(\cdot)$ is the marginal PDF of the given data $\rho$. We note that $\pi_{\text{obs}}(\rho)$ is generally unknown.
In our setting, the posterior given by (4) will depend nonlinearly on the random error $\delta$, due to using the approximate solution map (3) in the log-likelihood. Given a deterministic discretization error $\delta$, one can apply standard textbook results to bound the resulting error on the posterior (4) in, e.g., Hellinger distance [26, Ch. 1]. Similar results on posterior convergence were proven in [17] for random errors $\delta$, under suitable conditions, for the limit $\delta\to 0$. However, we consider the case of non-negligible stochastic errors. To the best of our knowledge, no theoretical results on well-posedness exist in this setting. We focus on algorithmic analysis in this work, rather than the posterior itself.
@@
In the considered setting with stochastic errors (3), we note that the use of MCMC for sampling the posterior (4) results in a nested sampling problem as each MCMC sample requires a full Monte Carlo simulation. Our interest lies in understanding the effect of non-negligible noise in the forward map. For inherently random but unbiased likelihoods, one can rely on pseudo-marginal MCMC methods [1]. Such likelihoods are often the result of a product or sum of individual contributions from different data points [19] or simulation outputs [28]. Pseudo-marginal methods consider the combined probability space of the parameter to be inferred and the stochastic dimension of the forward map. The aim is to approximate the marginal distribution of the inferred parameter $D$, i.e., integrating out the forward map’s stochasticity, using samples from this combined probability space. Although these methods are well studied, the assumption of unbiasedness means that they cannot be directly applied in our setting. We also note that similar work on nested sampling has been conducted in the setting of optimal experimental design, where nested Monte Carlo sampling is used to marginalize over so-called nuisance parameters [2, 8].
The goal of this work is to consider the non-asymptotic sampling regime in $P$, i.e., simulations with non-negligible and possibly biased stochastic error. We study how the Monte Carlo simulation error affects the ratio of computed approximate likelihood values, as a first step towards understanding the acceptance-rejection behavior of the Metropolis-Hastings algorithm in this setting. Notably, we observe that this ratio becomes a random variable, whose distribution we study in terms of its moments. We derive an explicit expression for these moments when they exist, which can inform practitioners in selecting a suitable value for $P$ that balances computational cost with accuracy on approximated acceptance rates. This line of work complements that presented in [20], applying MCMC to Monte Carlo simulations in photoacoustic imaging.
@@ 2 Monte Carlo simulation of the diffusion equation
In this Section, we derive a Monte Carlo discretization of the model (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) and comment on the structure of the resulting error. This derivation is purposefully brief. For a more detailed mathematical derivation of Monte Carlo simulation for parabolic PDEs, we refer to textbooks, such as [11]. We assume, for simplicity, that $\rho(x,t)\geq 0$. As such, we can interpret $\rho(x,t)$ as a probability density over $x$ at any given time $t$. We then make use of the fact that (LABEL:eq:diff_eq) is the Fokker-Planck equation to the stochastic differential equation (SDE)
@@
at time $t_{k}=k\Delta t$. Using an Euler-Maruyama discretization, we propagate each particle $p$ in (6) as a discretized realization of (5), i.e., for $p=1,\dots,P$
@@
and $X_{p,0}\overset{\text{i.i.d.}}{\sim}\rho_{0}$. Simulating such an ensemble, produces an ensemble of particles, where the values $X_{p,k}$ are distributed according to $\rho(x,t)$ for any $t_{k}=k\Delta t$.
@@
We define the error $\delta_{k}\in\mathbb{R}^{N}$ of the Monte Carlo simulation, at time $t_{k}=k\Delta t$, using the elementwise notation
$\delta_{n,k}=\rho(x_{n},t_{k})-\hat{\rho}_{n,k},\qquad n=0,\dots,N-1,\qquad k\geq 0.$ (9)
@@
$\mathbb{E}\left[\hat{\rho}_{n,k}\right]=\frac{1}{\Delta x}\int_{n\Delta x}^{(n+1)\Delta x}\rho(x,t_{k})\text{d}x=\frac{F((n+1)\Delta x,t_{k})-F(n\Delta x,t_{k})}{\Delta x},$ (11)
@@
$\mathbb{E}\left[\hat{\rho}_{n,k}\right]=\rho(x,t_{k})+\left(\left(n+\frac{1}{2}\right)\Delta x-x\right)\frac{\partial\rho}{\partial x}(x,t_{k})+\mathcal{O}(\Delta x^{2}).$ (12)
@@
$\left|\left(n+\frac{1}{2}\right)\Delta x-x\right|\leq\frac{\Delta x}{2}.$ (13)
@@
The finite ensemble size $P$ induces a sampling error that scales with the number of particles in a given cell. As each particle in the ensemble (6) is fully independent and produces a single contribution to the binning estimator (8), we observe by the central limit theorem that the variance of the individual grid cells scales with $\mathcal{O}\left(\frac{1}{P\Delta x}\right)$ as $P\Delta x\to\infty$. Moreover, the distribution of $\delta_{k}$ will converge to a multivariate Gaussian in this limit, with mean and variance corresponding to the bias and variance in the error decomposition (10). Hence, motivating the assumed distribution in (3). Despite the trajectory independence, we note that the grid cells themselves are correlated due to the fact that each trajectory can only score in one cell at a given time step.
@@ 3 Analysis of approximate likelihood ratios
is a random variable, whose $p$-th moment exists iff $p\Sigma_{\eta}\succ\Sigma(D_{2})$, with $\succ$ denoting the Löwner order.
@@
$\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\mathbb{R}^{N}$}}$}}\nolimits\!\!\exp\left(-\frac{\left\|\delta_{1}-\Delta\rho_{1}\right\|^{2}_{M^{-1}}}{2}\right)\frac{1}{\sqrt{|2\pi\Sigma_{1}|}}\exp\left(-\frac{\left\|\delta_{1}-\mu_{1}\right\|^{2}_{\Sigma_{1}^{-1}}}{2}\right)\text{d}\delta_{1},$ (19)
@@
$\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\mathbb{R}^{N}$}}$}}\nolimits\!\!\exp\left(-\frac{\left\|\delta_{1}-\alpha_{1}\right\|^{2}_{S_{1}^{-1}}}{2}\right)\frac{1}{\sqrt{|2\pi\Sigma_{1}|}}\exp\left(-C_{1}\right)\text{d}\delta_{1}$ (25)
$=\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\mathbb{R}^{N}$}}$}}\nolimits\!\!\exp\left(-\frac{\left\|\delta_{1}-\alpha_{1}\right\|^{2}_{S_{1}^{-1}}}{2}\right)\frac{1}{\sqrt{|2\pi\Sigma_{1}|}}\exp\left(-C_{1}\right)\text{d}\delta_{1}$ (26)
@@
$\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\mathbb{R}^{N}$}}$}}\nolimits\!\!\exp\left(\frac{\left\|\delta_{2}-\Delta\rho_{2}\right\|^{2}_{M^{-1}}}{2}\right)\frac{1}{\sqrt{|2\pi\Sigma_{2}|}}\exp\left(-\frac{\left\|\delta_{2}-\mu_{2}\right\|^{2}_{\Sigma_{2}^{-1}}}{2}\right)\text{d}\delta_{2}.$ (28)
@@
$\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\mathbb{R}^{N}$}}$}}\nolimits\!\!\exp\left(-\frac{\left\|\delta_{2,\perp}-\alpha_{2}\right\|^{2}_{S_{2}^{\dagger}}}{2}\right)\exp\left(\left\langle\delta_{2,0},\gamma_{2,0}\right\rangle\right)\frac{1}{\sqrt{|2\pi\Sigma_{2}|}}\exp\left(-C_{2}\right)\text{d}\delta_{2}.$ (34)
@@
Remark 3.4.
The proof of Theorem 3.1 is quite verbose, in order to derive explicit expressions for the $p$-th moment of the distribution of the ratio (15). One can follow a much shorter argument to state that there exists an upper bound for $\Sigma(D_{2})$ as a function of the number of desired moments, for example by observing that the numerator of (15) is bounded and applying Fernique’s theorem [5] to the denominator.
@@ 4 Numerical experiments
We now perform a numerical experiment to confirm the observations in Theorem 3.1. To this end, we use the model problem (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) and consider transitions between fixed values for $D_{1}$ and $D_{2}$. We produce synthetic observations, using a finite difference discretization whose solution is perturbed with noise distributed according to (2). Throughout this section we set $N=100$, $L=10$, $t=10$ and $D^{\ast}=0.1$. The reference solution to produce the observation is computed with $\Delta t=0.1$, while the Monte Carlo simulations use $\Delta t=t$ to cut on computational cost, thanks to the unbiasedness of (7). For $D_{1}$ and $D_{2}$, we consider differing combinations of these parameters from the set $\{0.08,0.1\}$, i.e., the true value and a similar, but less probable, value. The code used to generate the results in this section can be found at github.com/UQatKIT/FrontUQ-2024-proceedings-likelihood-ratios.
@@
Figure 1: Computing the expectation of the approximate likelihoods and the ratio of these expectations, given a discrete solution of the problem (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) perturbed by synthetic observation noise, for $D^{\ast}=D_{1}=0.1$ and $D_{2}=0.08$.
@@
Figure 2: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) perturbed by synthetic observation noise for $D^{\ast}=D_{1}=D_{2}=0.1$.
@@
Figure 3: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) perturbed by synthetic observation noise for $D^{\ast}=D_{2}=0.1$ and $D_{1}=0.08$.
@@
Figure 4: Computing the expectation of the approximate likelihood ratio given a discrete solution of the problem (LABEL:eq:diff_eq)–(LABEL:eq:diff_eq_boundary) perturbed by synthetic observation noise for $D^{\ast}=D_{1}=0.1$ and $D_{2}=0.08$.
Next, in Figure 2, we consider the case where $D_{2}=D_{1}=D^{\ast}=0.1$, i.e., both the current state and proposal match the true parameter value. We plot the expectation of the approximate likelihood ratio in Figure 2(a). Here we observe a significant blowup in the expectation at approximately $\sigma_{\eta}=\sigma_{\delta}$, as predicted by Theorem 3.1. For sufficiently small values of $\sigma_{\delta}$, we observe that the expectation of the likelihood ratio converges to the true value of 1. To better understand the divergence of the approximate likelihood ratio, we plot the expectation of the minimum of the ratio and 1 in Figure 2(b), which in the case of $D_{1}=D_{2}$ directly corresponds to the acceptance probability in the Metropolis-Hastings algorithm, see e.g. [25, Ch. 7]. Here, it becomes clear that when the approximate likelihood ratio diverges, the acceptance probability converges to the value 0.5. This behavior gives insight into the precise mechanism behind the divergence. Namely, in this regime, the stochastic error in the approximate likelihood evaluations is sufficiently large that the ratio consists of a numerator and denominator with wildly differing orders of magnitude. Hence, the random variable converges to a distribution with two states, one of which diverges to infinity, the other converges to zero. Each state has probability 0.5.